A 2.55 inch diameter pipe is required to carry 46.1 gpm. What flow velocity is required, in ft/s?
Answer #1

Q = 46.4 gpm = (46.1) (0.003785 MB) / 60 0 - 2.908 x109 miles Q = Ar Q = (161) (0.1347+²) 1608 Q = 0.1027 ft3/ Q = An 0-1027